OpenStudy (anonymous):

how to solve ((x+1)-1) 2(squared)

OpenStudy (anonymous):

(x+1)^2 -2 (1)(x+1)+(1)^2 x^2+2x+1-2x-2+1 x^2

OpenStudy (anonymous):

the answer is x(squared) but i dont know how??

OpenStudy (anonymous):

or just simplify (x+1-1)=x and square it

OpenStudy (anonymous):

very confusing

OpenStudy (anonymous):

(a-b)^2 = a^2 -2ab + b^2 a=(x+1) b=1

OpenStudy (anonymous):

tnx i think i got it ahaha

OpenStudy (anonymous):

sir that is a negative 1 or -1 how it becomes a positive 1??

OpenStudy (anonymous):

it is either a+b where b=-1 and then it is a^2+2ab+b^2 or a-b where b=1 and a^2-2ab+b^2 it comes out the same either way

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

x^{2}